Ms X drove at the speed of 45 Km/hr from home to a resort. Returning over the same route she got stuck in traffic and took an hour longer, also she could drive only at the speed of 40 Km/hr. How many kilometers did she drive each way?
Answer by mananth(16949)

let resort distance be x miles
x/45 = time to go to resort.
...
return
time = x/40
...
x/40-x/45=1
..
45x-40x /40*45 =1
5x=1800
/5
x=360 miles
...
m.ananth@hotmail.ca